There are 6.02 x 10 23 particles in 22.4 liters of air. Calculate the number of particles in 10L of air.




1
Expert's answer
2021-02-03T07:06:32-0500

6.02*1023 ---------- 22.4 L

N ---------------- 10 L

N = 10*6.02*1023/22.4 = 2.6875*1023

Answer: 2.69*1023 particles